OpenStudy (gorica):

Do you know where can I find how to get relations between heat capacities Cp and Cv for ideal and real gases? Or someone can write it here?

OpenStudy (anonymous):

You can always start with Cv=3/2R for a monatomic particle and use Cp=Cv+R. Cv= 5/2R for Diatomic CV= 3R for Polyatomic R is a constant that you know right? Cheers :D

OpenStudy (gorica):

I have to show why is it Cp=Cv+nR for ideal gas and why it is what it is for real gas. I found this, and I am ok with ideal gas... but I am not really clear with real gas :/

OpenStudy (anonymous):

OooOOooo, That complicates things then :( I'm not sure I can help you further, sorry Cheers :D

OpenStudy (gorica):

thanks anyway :)

OpenStudy (ghazi):

hope this will help you out , hold a sec

OpenStudy (ghazi):

if you are not able to read it , let me know :)

OpenStudy (gorica):

I think I found what I was looking for... via coefficient of volume expansion Thank you very much

OpenStudy (ghazi):

\[h= u+PV\]\[dh=du+d(PV)\] \[dh=du+d(RT), dh=du+RdT\]\[\frac{ dh }{ dT }=\frac{ du }{ dT }+R\]\[C _{p}=C _{v}+R, C _{p}-C _{v}=R\] you can use this derivation too

OpenStudy (ghazi):

and you are welcome :)

OpenStudy (gorica):

what is h?

OpenStudy (ghazi):

h-- enthalpy u -- internal energy dh/dT -- Cp du/dT-- Cv

OpenStudy (gorica):

is Cp defines as dh/dT or we get it somehow because p=const?

OpenStudy (gorica):

*defined

OpenStudy (ghazi):

yep, Cp is specific heat capacity at constant pressure, and change in enthalpy at with temperature is the same quantity , for ideal gas dh=CpdT and gamma = Cp/Cv

OpenStudy (gorica):

ok. I will read this all again and if I have more questions, I will write them tomorrow, ok? :) thank you again :)

OpenStudy (ghazi):

sure :)

OpenStudy (gorica):

do you maybe have graphs of isotherms of Van der Waals gas and real isotherms and something about thermometer with thermocouple, I can't find it on internet?

OpenStudy (gorica):

:) thanks :)